Aurora Co. has beginning inventory of 15,000 units at a cost of $34,000. Total sales of 225,000 units were $865,000. The follQuestions:

1. Ending Inventory?

2. Total cost available for sale?

3. Total number of units available for sale?

4. Ending inventory if the company uses LIFO?

5. Gross profit if the company uses FIFO?

6. Cost of goods sold if the company uses average cost?

7. Ending inventory if the company uses average cost?

8. gross profit if the company uses average cost?
Answer #1

FIFO = THE GOODS PURCHASED EARLIER ARE SOLD FIRST

lifo= the goods latest purchased are sold

average = average cost is found out to determine cost of goods sold

1. Ending Inventory?

=BEGINNING +PURCHASE-SALES

=15000+[60000+50000+50000+70000]-225000

=20000 UNITS

2.Total cost available for sale?

BEGINNING+PURCHASE

=34000+144000+125000+130000+192500

=625500$

3.Total number of units available for sale?

=15000+60000+50000+50000+70000

=245000 U NITS

5. FIFO

COST OF GOODS SOLD 570500 [34000+60000*2.4+50000*2.5+50000*2.6+50000*2.75]
ENDING INVENTORY 55000 [20000*2.75]]
GROSS PROFIT 294500 [865000-570500]SALES-COST OF GOODS SOLD

4.Ending inventory if the company uses LIFO?

COST OF GOODS SOLD 579500 [70000*2.75+50000*2.6+50000*2.5+55000*2.4]
ENDING INVENTORY 46000 [5000*2.4+34000]
GROSS PROFIT 285500 [865000-579500]SALES-COST OF GOODS SOLD

5.AVERAGE

BEGINNING 15000 34000
PURCHASE 60000 2.4 144000
50000 2.5 125000
50000 2.6 130000
70000 2.75 192500
TOTAL UNITS 245000 625500

AVERAGE = 625500/245000=2.55$

COST OF GOODS SOLD = 225000*2.55

=573750$

ENDING INVENTORY = 2.55$*20000UNITS

=51000$

GROSS PROFIT =865000-573750

=291250$